if a bus travels 300km at a given speed. if the speed is increased to 10 kph the trip would have been reduced by one hour.find the speed of the bus.
Answer by mananth(16949)

let the speed of the bus be x
increased by 10
so speed = x+10
time with original speed - time with new speed = 1
300/x-300/(x+10)=1
LCD = x(x+10)
300*(x+10)-300x=x(x+10)
300x+3000-300x=x^2+10x
x^2+10x-3000=0
x^2+60x-50x-3000=0
x(x+60)-50(x+60)=0
(x-50)(x+60)=0
x=50 which is positive, the speed of the bus